# MyFramelinecolor=black, outerlinewidth=2pt, innertopmargin=4pt, innerbottommargin=4pt, innerrightmargin=4pt, innerleftmargin=4pt, leftmargin = 4pt, rightmargin = 4pt, backgroundcolor=gray!40

### v1

@@ 6 Limitations
DistR’s design has three limitations. First, although DistR permits the use of unsafe code, its consistency guarantees are only applicable to safe Rust code. In unsafe code blocks, developers are responsible for ensuring consistency themselves. Second, DistR’s superior performance relies on SWMR semantics exposed by applications. In cases where data is mostly under shared states (such as Mutex), DistR degenerates into a traditional DSM system; all concurrent accesses to the same data have to be centralized and serialized by the server responsible for the shared states. However, such scenarios contradict Rust’s recommended programming practices. Finally, the current implementation of DistR does not support address space layout randomization (ASLR) yet, and we have temporarily disabled it. However, DistR’s design is compatible with ASLR as long as DistR threads share the same randomized address space layout on each server. This can be achieved by delegating the randomization of stack and heap address allocation in DistR to its global controller, a feature that will be supported in future versions of DistR.
@@ 7 Evaluation
Methodology. We compared DistR with two state-of-the-art DSM systems, GAM [14] and Grappa [60]. For a fair comparison, we ported the evaluated applications to each baseline system and invested extensive effort in tuning parameters to achieve their best possible performance. GAM offers ordinary object read/write interfaces, and we exported it as a library to Rust and hooked pointer dereferencing to use GAM’s API without program modification. Grappa, in contrast, offers a drastically different programming abstraction that requires rewriting the program to access shared memory via delegation. Therefore, we re-implemented applications in C++ and re-structured them using Grappa’s abstractions.
@@ 7.1 Applications
Figure 5: Application throughput when running with DistR, GAM, and Grappa, normalized to the throughput of their original implementation running on a single node. The number in the parenthesis is the original application’s throughput on a single node.
@@
SocialNet is a twitter-like web service from the DeathStarBench suite [33]. It is composed of 12 microservices with complicated call dependencies. Each microservice in SocialNet can scale independently with replicas, thereby offering higher throughput with more servers. SocialNet decouples the process of user texts, media resources, and storage into different microservices, and it employs RPCs to pass values (texts, media files, etc.) between them. DistR enables SocialNet to pass only references in RPCs, eliminating the serialization/deserialization overhead and redundant data copies. Because SocialNet was implemented in C++ and deployed with Docker Swarm [29], we ported it into Rust for our evaluation. We followed its original microservice structure but changed the RPC call sites to pass references instead of values, and we followed the original orchestration configuration to spread and scale each microservice in the cluster. We did not use any affinity annotations for SocialNet.
@@ 7.2 Scaling Performance
In this experiment, we investigated whether DistR can speed up applications by distributing them in a cluster and how well they can scale with the number of servers used. For each application, we first ran it as is on a single server without using DSM and measured its throughput. Then, we ran the same application on DSM (subject to modifications when running Grappa) with the same configuration but on varying numbers of servers and measured the throughput normalized to its single-node throughput (i.e., strong scaling). As GAM and Grappa cannot adaptively balance the workload across servers, we evenly distributed the application’s working set and threads among all participating nodes. Ideally, an application should scale linearly and enjoy proportionally higher throughput with more nodes. However, this is usually unachievable because of the limited parallelism of real-world applications and the coherence overhead of DSM systems, and a good result for DistR will show that applications’ throughput can get close to their ideal throughput.
Figure 5 shows the results for each application respectively. DistR outperforms both GAM and Grappa in all cases. On a single node, it is 1.04–2.10$\times$ faster than two baseline DSMs, while only adding a maximum overhead of 2.42% compared to the original program. When running with multiple nodes, DistR scales up applications significantly better than GAM and Grappa. On eight nodes, DistR achieves a throughput that is 1.33–2.64$\times$ higher than that of GAM, 2.53–29.16$\times$ higher than that of Grappa.
Compared to each program’s single-machine performance, using DSM over DistR enables each program to easily leverage the available distributed resources and achieve a throughput that is 3.34–11.73$\times$ higher than their single-machine counterparts. Next, we discuss each application to explain the scalability difference between DistR and the baseline DSMs.
DataFrame. As shown in Figure 5a, compared with its original version, DataFrame running on eight nodes with DistR achieves 5.57$\times$ higher throughput, whereas with GAM and Grappa, the throughput improvements are 2.18$\times$ and 1.69$\times$, respectively. In other words, DataFrame with DistR is 2.56$\times$ and 3.29$\times$ faster than GAM and Grappa on eight nodes, respectively.
@@
DistR outperforms GAM and scales much better because of its light coherence protocol, which incurs negligible object move overhead for writes and no coherence overhead for reads. The use of affinity annotations also helps DataFrame colocate worker threads with their frequently accessed data, bringing 20% additional boost (details in §7.3). GAM, in contrast, has to invalidate each cache block upon each write and read, thereby bottlenecked by the extensive coherence traffic. Grappa performs the worst in all three DSM systems due to its always-delegation programming model, which implements every global memory read/write via a delegated function call. The cost for delegation overwhelms the actual memory access latency in this case, ruining the performance of the shared hash table. Grappa’s delegation overhead actually causes a 1.23$\times$ slowdown when scaling DataFrame from a single node to two nodes.
@@
SocialNet. Since SocialNet is microservice-based and can be deployed distributively, we added another baseline by running the original (non-DSM) code but deploying it on varying numbers of nodes. Figure 5b demonstrates the performance of all systems. SocialNet runs consistently faster with all three DSM systems compared to the original version. DistR, GAM, and Grappa achieve a 2.18$\times$, 2.02$\times$, and 1.57$\times$ speedup on a single node and a 3.51$\times$, 1.33$\times$, and 1.39$\times$ speedup on eight nodes, respectively. In the conventional setup, SocialNet requires data—such as text and media files—to be serialized into byte streams for network transmission, and then deserialized back into usable formats at the receiving end. This serialization and deserialization process is computationally intensive, particularly for large or complex data objects. In contrast, DSM systems enable SocialNet to pass references instead of the entire data values required by remote procedure calls. This approach eliminates the need for serialization and deserialization, reduces redundant data copies, and significantly enhances performance. DistR scales much better than GAM and Grappa thanks to its lightweight coherence protocol, achieving up to 2.77$\times$ and 3.16$\times$ higher throughput than GAM and Grappa, respectively.
@@
GEMM. GEMM differs from the previous two applications in its high compute intensity and relatively infrequent shared memory accesses. In this application, matrices are transformed and divided into smaller sub-matrices for parallel processing. Each computing thread, responsible for multiplying sub-matrices, is assigned to a server. These threads cache their respective sub-matrices in the server’s local memory and access them repeatedly to compute product results. This process is highly compute-intensive. As depicted in Figure 5c, DistR and GAM scale well for GEMM and achieve 5.93$\times$, 3.82$\times$ speedup with eight nodes. In contrast, Grappa only achieves a 2.02$\times$ speedup with eight nodes due to its inability to cache sub-matrices locally, necessitating frequent remote accesses. DistR’s superior performance over GAM, with a 1.55$\times$ higher speedup on eight nodes, is primarily due to its more efficient handling of initial cross-server data accesses required when a sub-matrix is first accessed remotely. Unlike GAM, which incurs significant runtime overhead due to the maintenance of state and location of data copies, DistR directly copies data to local memory, without any complex cross-server synchronization operations, thus enhancing overall efficiency.
@@
Figure 5d shows the results. KV Store experiences a slowdown on all three DSM systems when scaling from a single node to two nodes (13% for DistR, 25% for GAM, and 93% for Grappa). However, the impact is mitigated when more servers are enlisted—DistR and GAM achieve 3.34$\times$ and 2.50$\times$ higher throughput on eight nodes compared to the original KV Store implementation, respectively. Due to the limited ownership semantics exposed by mutexes, DistR does not scale as well with KV Store as with other applications. DistR is 1.33$\times$ faster than GAM on eight nodes, benefiting from its adaptive load balancing and a more efficient implementation of mutexes utilizing one-sided RDMA atomic verbs, whereas GAM depends on less efficient two-sided RDMA messages for synchronization. Grappa, in contrast, incurs the highest distribution overhead and poorest scalability, primarily because each PUT/GET operation requires remote delegation, and nodes handling popular objects become bottlenecked due to skewed load.
@@ 7.3 Drill-Down Experiments
Figure 6: Effectiveness of DistR’s affinity annotations.
Latency (cycles) Average Median P90 DistR 395 356 536 Rust 364 332 496
Table 2: DistR’s Box pointer only adds a small dereferencing cost compared to Rust’s ordinary Box.
Affinity Annotations. In this experiment, we evaluated the individual contributions of affinity annotations by enabling each of them incrementally for DataFrame on eight nodes. Figure 6 reports the results. Using TBox helps DataFrame group chunks from the same column and eliminates the runtime dereference check overhead for single-column operations (e.g., filter), bringing a 12% throughput improvement. Adding spawn_to further improves the throughput by 9% by informing DistR runtime to colocate the worker thread to its input columns, which reduces cross-server memory accesses.
Runtime Dereference Checks. We measured the latency of dereferencing DistR’s Box pointer and compared it with an ordinary Rust Box pointer. Both of them point to an 8-byte object in local memory and not in CPU’s cache, which represents the common path for pointer dereferencing. Table 2 reports the results. DistR only adds a small overhead of $\sim$30 cycles. Note that this microbenchmark is extremely memory-intensive, whereas real-world applications usually employ larger object sizes and are more compute-intensive, further mitigating the runtime check overhead. For our evaluated applications, we observed a 1.02% overhead for DataFrame and a 1.14% overhead for BLAS, when they run with DistR on a single node, respectively.
@@
Figure 7: Comparison of cache coherence costs between DistR, GAM, and Grappa on eight nodes.
Cost of Cache Coherence. In this experiment, we ran each application again on a single node and eight nodes but fixed the total amount of CPU and memory resources. For the eight-node setting, we distributed the resources evenly to each node and measured application throughput. We expect to see a slowdown due to the cost of running the coherence protocol and cross-server memory accesses, but a good result for DistR should show that application performance remains close to its original single-node version. Figure 7 reports the results. SocialNet uses pass-by-value RPCs in its original version and is significantly slower than our DSM-based version, so it is omitted in the evaluation. DistR adds only moderate cache coherence cost with an overhead of 32% in the worst case (KV Store) and 4% in the best case (GEMM). GAM and Grappa, in contrast, incur much higher overheads ranging from 10% to 98% for different applications.
@@ 8 Related Work
Disaggregated and Remote Memory. Memory disaggregation and remote memory techniques are another promising approach to scaling applications out of a single machine. Their key idea is to connect a host server with large memory pools [34, 47, 41] via fast datacenter network, which can be accessed by applications via OS kernel [76, 4, 86, 69] or software runtimes [73, 84, 85, 53, 35]. However, they do not provide cache coherence.
Distributed Programming Abstractions. Researchers have studied and proposed new programming languages and abstractions. Munin[10] built a type system that defines types for local and global pointers and tracks whether the pointer is shared via type checking. X10 [20, 40] and UPC [31] introduce function offloading interfaces for distributed computing and additional type annotations to reduce the runtime overhead. Ray [88] and Nu [72] are two recent systems proposing new abstractions for distributed programming. Unlike DistR, they require effort to port applications to avoid fine-grained memory sharing.
Hardware-Accelerated DSM. Specialized datacenter network technologies and emerging hardware designs stand for another trend to accelerate DSM. Clio [36], StRoM [79], and RMC [5] reduce remote memory access latency by offloading tasks into customized hardware. Concordia [87], Kona [15], and CXL [23, 24, 26, 25, 27, 91, 49, 48] enable block-level or cache-line-level memory coherence with their hardware-implemented protocols. DistR can benefit from advances in hardware support and achieve better scalability.
@@ 9 Conclusion
This paper presents DistR, a practical DSM system based on the ownership model. It automatically turns a single-machine Rust program into its distributed version with a lightweight coherence protocol guided by language semantics. DistR significantly outperforms existing state-of-the-art DSM systems, demonstrating that a language-guided DSM can achieve strong memory consistency, transparency, and efficiency simultaneously.
@@ A.1 Artifact Summary
DistR is an efficient, consistent, and user-friendly DSM system featuring a lightweight coherence protocol guided by language semantics. DistR allows for seamless scaling of single-machine applications to multi-server environments without sacrificing performance. Demonstrating significant improvements over existing DSM systems, DistR combines strong memory consistency, transparency, and efficiency effectively.
@@ A.3 Description
A.3.1 DistR’s Codebase
DistR comprises four main components:
@@ A.3.1 DistR’s Codebase
    Applications integrated with DistR
@@
A.3.2 Deploying DistR
The initial step in deploying DistR involves cloning the source code on all involved servers: {mdframed}[style=MyFrame,nobreak=true] git clone git@github.com:uclasystem/DRust.git

### v2

@@ 6 Limitations
DRust’s design has three limitations. First, although DRust permits the use of unsafe code, its consistency guarantees are only applicable to safe Rust code. In unsafe code blocks, developers are responsible for ensuring consistency themselves. Second, DRust’s superior performance relies on SWMR semantics exposed by applications. In cases where data is mostly under shared states (such as Mutex), DRust degenerates into a traditional DSM system; all concurrent accesses to the same data have to be centralized and serialized by the server responsible for the shared states. However, such scenarios contradict Rust’s recommended programming practices. Finally, the current implementation of DRust does not support address space layout randomization (ASLR) yet, and we have temporarily disabled it. However, DRust’s design is compatible with ASLR as long as DRust threads share the same randomized address space layout on each server. This can be achieved by delegating the randomization of stack and heap address allocation in DRust to its global controller, a feature that will be supported in future versions of DRust.
@@ 7 Evaluation
Methodology. We compared DRust with two state-of-the-art DSM systems, GAM [14] and Grappa [60]. For a fair comparison, we ported the evaluated applications to each baseline system and invested extensive effort in tuning parameters to achieve their best possible performance. GAM offers ordinary object read/write interfaces, and we exported it as a library to Rust and hooked pointer dereferencing to use GAM’s API without program modification. Grappa, in contrast, offers a drastically different programming abstraction that requires rewriting the program to access shared memory via delegation. Therefore, we re-implemented applications in C++ and re-structured them using Grappa’s abstractions.
@@ 7.1 Applications
Figure 5: Application throughput when running with DRust, GAM, and Grappa, normalized to the throughput of their original implementation running on a single node. The number in the parenthesis is the original application’s throughput on a single node.
@@
SocialNet is a twitter-like web service from the DeathStarBench suite [32]. It is composed of 12 microservices with complicated call dependencies. Each microservice in SocialNet can scale independently with replicas, thereby offering higher throughput with more servers. SocialNet decouples the process of user texts, media resources, and storage into different microservices, and it employs RPCs to pass values (texts, media files, etc.) between them. DRust enables SocialNet to pass only references in RPCs, eliminating the serialization/deserialization overhead and redundant data copies. Because SocialNet was implemented in C++ and deployed with Docker Swarm [28], we ported it into Rust for our evaluation. We followed its original microservice structure but changed the RPC call sites to pass references instead of values, and we followed the original orchestration configuration to spread and scale each microservice in the cluster. We did not use any affinity annotations for SocialNet.
@@ 7.2 Scaling Performance
In this experiment, we investigated whether DRust can speed up applications by distributing them in a cluster and how well they can scale with the number of servers used. For each application, we first ran it as is on a single server without using DSM and measured its throughput. Then, we ran the same application on DSM (subject to modifications when running Grappa) with the same configuration but on varying numbers of servers and measured the throughput normalized to its single-node throughput (i.e., strong scaling). As GAM and Grappa cannot adaptively balance the workload across servers, we evenly distributed the application’s working set and threads among all participating nodes. Ideally, an application should scale linearly and enjoy proportionally higher throughput with more nodes. However, this is usually unachievable because of the limited parallelism of real-world applications and the coherence overhead of DSM systems, and a good result for DRust will show that applications’ throughput can get close to their ideal throughput.
Figure 5 shows the results for each application respectively. DRust outperforms both GAM and Grappa in all cases. On a single node, it is 1.04–2.10$\times$ faster than two baseline DSMs, while only adding a maximum overhead of 2.42% compared to the original program. When running with multiple nodes, DRust scales up applications significantly better than GAM and Grappa. On eight nodes, DRust achieves a throughput that is 1.33–2.64$\times$ higher than that of GAM, 2.53–29.16$\times$ higher than that of Grappa.
Compared to each program’s single-machine performance, using DSM over DRust enables each program to easily leverage the available distributed resources and achieve a throughput that is 3.34–11.73$\times$ higher than their single-machine counterparts. Next, we discuss each application to explain the scalability difference between DRust and the baseline DSMs.
DataFrame. As shown in Figure 5a, compared with its original version, DataFrame running on eight nodes with DRust achieves 5.57$\times$ higher throughput, whereas with GAM and Grappa, the throughput improvements are 2.18$\times$ and 1.69$\times$, respectively. In other words, DataFrame with DRust is 2.56$\times$ and 3.29$\times$ faster than GAM and Grappa on eight nodes, respectively.
@@
DRust outperforms GAM and scales much better because of its light coherence protocol, which incurs negligible object move overhead for writes and no coherence overhead for reads. The use of affinity annotations also helps DataFrame colocate worker threads with their frequently accessed data, bringing 20% additional boost (details in §7.3). GAM, in contrast, has to invalidate each cache block upon each write and read, thereby bottlenecked by the extensive coherence traffic. Grappa performs the worst in all three DSM systems due to its always-delegation programming model, which implements every global memory read/write via a delegated function call. The cost for delegation overwhelms the actual memory access latency in this case, ruining the performance of the shared hash table. Grappa’s delegation overhead actually causes a 1.23$\times$ slowdown when scaling DataFrame from a single node to two nodes.
@@
SocialNet. Since SocialNet is microservice-based and can be deployed distributively, we added another baseline by running the original (non-DSM) code but deploying it on varying numbers of nodes. Figure 5b demonstrates the performance of all systems. SocialNet runs consistently faster with all three DSM systems compared to the original version. DRust, GAM, and Grappa achieve a 2.18$\times$, 2.02$\times$, and 1.57$\times$ speedup on a single node and a 3.51$\times$, 1.33$\times$, and 1.39$\times$ speedup on eight nodes, respectively. In the conventional setup, SocialNet requires data—such as text and media files—to be serialized into byte streams for network transmission, and then deserialized back into usable formats at the receiving end. This serialization and deserialization process is computationally intensive, particularly for large or complex data objects. In contrast, DSM systems enable SocialNet to pass references instead of the entire data values required by remote procedure calls. This approach eliminates the need for serialization and deserialization, reduces redundant data copies, and significantly enhances performance. DRust scales much better than GAM and Grappa thanks to its lightweight coherence protocol, achieving up to 2.77$\times$ and 3.16$\times$ higher throughput than GAM and Grappa, respectively.
@@
GEMM. GEMM differs from the previous two applications in its high compute intensity and relatively infrequent shared memory accesses. In this application, matrices are transformed and divided into smaller sub-matrices for parallel processing. Each computing thread, responsible for multiplying sub-matrices, is assigned to a server. These threads cache their respective sub-matrices in the server’s local memory and access them repeatedly to compute product results. This process is highly compute-intensive. As depicted in Figure 5c, DRust and GAM scale well for GEMM and achieve 5.93$\times$, 3.82$\times$ speedup with eight nodes. In contrast, Grappa only achieves a 2.02$\times$ speedup with eight nodes due to its inability to cache sub-matrices locally, necessitating frequent remote accesses. DRust’s superior performance over GAM, with a 1.55$\times$ higher speedup on eight nodes, is primarily due to its more efficient handling of initial cross-server data accesses required when a sub-matrix is first accessed remotely. Unlike GAM, which incurs significant runtime overhead due to the maintenance of state and location of data copies, DRust directly copies data to local memory, without any complex cross-server synchronization operations, thus enhancing overall efficiency.
@@
Figure 5d shows the results. KV Store experiences a slowdown on all three DSM systems when scaling from a single node to two nodes (13% for DRust, 25% for GAM, and 93% for Grappa). However, the impact is mitigated when more servers are enlisted—DRust and GAM achieve 3.34$\times$ and 2.50$\times$ higher throughput on eight nodes compared to the original KV Store implementation, respectively. Due to the limited ownership semantics exposed by mutexes, DRust does not scale as well with KV Store as with other applications. DRust is 1.33$\times$ faster than GAM on eight nodes, benefiting from its adaptive load balancing and a more efficient implementation of mutexes utilizing one-sided RDMA atomic verbs, whereas GAM depends on less efficient two-sided RDMA messages for synchronization. Grappa, in contrast, incurs the highest distribution overhead and poorest scalability, primarily because each PUT/GET operation requires remote delegation, and nodes handling popular objects become bottlenecked due to skewed load.
@@ 7.3 Drill-Down Experiments
Figure 6: Effectiveness of DRust’s affinity annotations.
Latency (cycles) Average Median P90 DRust 395 356 536 Rust 364 332 496
Table 2: DRust’s Box pointer only adds a small dereferencing cost compared to Rust’s ordinary Box.
Affinity Annotations. In this experiment, we evaluated the individual contributions of affinity annotations by enabling each of them incrementally for DataFrame on eight nodes. Figure 6 reports the results. Using TBox helps DataFrame group chunks from the same column and eliminates the runtime dereference check overhead for single-column operations (e.g., filter), bringing a 12% throughput improvement. Adding spawn_to further improves the throughput by 9% by informing DRust runtime to colocate the worker thread to its input columns, which reduces cross-server memory accesses.
Runtime Dereference Checks. We measured the latency of dereferencing DRust’s Box pointer and compared it with an ordinary Rust Box pointer. Both of them point to an 8-byte object in local memory and not in CPU’s cache, which represents the common path for pointer dereferencing. Table 2 reports the results. DRust only adds a small overhead of $\sim$30 cycles. Note that this microbenchmark is extremely memory-intensive, whereas real-world applications usually employ larger object sizes and are more compute-intensive, further mitigating the runtime check overhead. For our evaluated applications, we observed a 1.02% overhead for DataFrame and a 1.14% overhead for BLAS, when they run with DRust on a single node, respectively.
@@
Figure 7: Comparison of cache coherence costs between DRust, GAM, and Grappa on eight nodes.
Cost of Cache Coherence. In this experiment, we ran each application again on a single node and eight nodes but fixed the total amount of CPU and memory resources. For the eight-node setting, we distributed the resources evenly to each node and measured application throughput. We expect to see a slowdown due to the cost of running the coherence protocol and cross-server memory accesses, but a good result for DRust should show that application performance remains close to its original single-node version. Figure 7 reports the results. SocialNet uses pass-by-value RPCs in its original version and is significantly slower than our DSM-based version, so it is omitted in the evaluation. DRust adds only moderate cache coherence cost with an overhead of 32% in the worst case (KV Store) and 4% in the best case (GEMM). GAM and Grappa, in contrast, incur much higher overheads ranging from 10% to 98% for different applications.
@@ 8 Related Work
Disaggregated and Remote Memory. Memory disaggregation and remote memory techniques are another promising approach to scaling applications out of a single machine. Their key idea is to connect a host server with large memory pools [33, 46, 40] via fast datacenter network, which can be accessed by applications via OS kernel [76, 4, 86, 69] or software runtimes [73, 84, 85, 52, 34, 87]. However, they do not provide cache coherence.
Distributed Programming Abstractions. Researchers have studied and proposed new programming languages and abstractions. Munin[10] built a type system that defines types for local and global pointers and tracks whether the pointer is shared via type checking. X10 [20, 39] and UPC [30] introduce function offloading interfaces for distributed computing and additional type annotations to reduce the runtime overhead. Ray [89] and Nu [72] are two recent systems proposing new abstractions for distributed programming. Unlike DRust, they require effort to port applications to avoid fine-grained memory sharing.
Hardware-Accelerated DSM. Specialized datacenter network technologies and emerging hardware designs stand for another trend to accelerate DSM. Clio [35], StRoM [79], and RMC [5] reduce remote memory access latency by offloading tasks into customized hardware. Concordia [88], Kona [15], and CXL [23, 24, 26, 25, 92, 48, 47] enable block-level or cache-line-level memory coherence with their hardware-implemented protocols. DRust can benefit from advances in hardware support and achieve better scalability.
@@ 9 Conclusion
This paper presents DRust, a practical DSM system based on the ownership model. It automatically turns a single-machine Rust program into its distributed version with a lightweight coherence protocol guided by language semantics. DRust significantly outperforms existing state-of-the-art DSM systems, demonstrating that a language-guided DSM can achieve strong memory consistency, transparency, and efficiency simultaneously.
@@ A.1 Artifact Summary
DRust is an efficient, consistent, and user-friendly DSM system featuring a lightweight coherence protocol guided by language semantics. DRust allows for seamless scaling of single-machine applications to multi-server environments without sacrificing performance. Demonstrating significant improvements over existing DSM systems, DRust combines strong memory consistency, transparency, and efficiency effectively.
@@ A.3 Description
A.3.1 DRust’s Codebase
DRust comprises four main components:
@@ A.3.1 DRust’s Codebase
    Applications integrated with DRust
@@
A.3.2 Deploying DRust
The initial step in deploying DRust involves cloning the source code on all involved servers: {mdframed}[style=MyFrame,nobreak=true] git clone git@github.com:uclasystem/DRust.git